OpenStudy (anonymous):

Solve 5x^2 = –30x – 65 x = –3 ± 2i x = –3 ± 4i x = –6 ± 2i x = –6 ± 4i

OpenStudy (aroub):

x = –3 ± 2i

OpenStudy (anonymous):

5x^2 + 30x +65 =0 x^2 + 6x + 13 = 0 (x+3)^2 +4 = 0 x+3 = sqrt(-4) = plus minus 2i x = –3 ± 2i